# Generative AI in Transportation Planning: A Survey

### v1

@@ 5.8 Case Study: LLM’s Ability in OD-Calibration in Terms of Feature Quality
This subsection presents a case study on how LLMs perform in Origin-Destination (OD) Calibration tasks for transportation planning. Specifically, we utilize an O-D calibration dataset from Open Traffic Intelligence (OpenTI) [7], which includes initial OD matrices and observed traffic counts, to evaluate how generative AI models perform when tasked with OD calibration. The study focuses on two critical aspects: (1) Exploring the few-shot learning solution designs in current LLMs for solving the O-D calibration task, and (2) assessing the quality of model-generated features (observation counts) compared to ground-truth OD matrix’s count feature, an overview of the problem is as shown in the Figure 6.
@@ 5.8.1 Task Description
In this section, we consider an origin-destination (O-D) matrix $\mathbf{X}$ of dimension $T\times 56\times 56$, where $T$ is the total number of discrete time slices and $56$ represents the number of blocks in the city. Specifically, $\mathbf{X}$ contains the entries $x_{t,i,j}$ for $t=1,\dots,T$ and $i,j\in\{1,\dots,56\}$, where $x_{t,i,j}\geq 0$ denotes the traffic demand (number of vehicles) from block $i$ to block $j$ during time slice $t$. Once the full matrix $\mathbf{X}$ is specified, a traffic simulator [280] is executed to generate network flows; however, since the observation point is sparse, the effect from the $f\bigl{(}\mathbf{X}\bigr{)}$ is very limited, which bring much difficulty to this task. If we let $O_{t}$ be the ground-truth measurement at the sensor for time slice $t$. Our goal is to calibrate the unknown entries of $\mathbf{X}$ so as to minimize the mean squared error between the simulator-derived flows and the real observations. Concretely, we define the following:
@@ 5.8.2 Model Configurations, Computational Resources, and Dataset Selection
We conduct exploration on the open-sourced and widely-used generative AI models Llama 3.1-70B [281]—to simulate OD calibration scenarios. Experiments are conducted on hardware configurations tailored for large-scale traffic data tasks and are deployed on a server with 8 NVIDIA A100 GPUs, 2 AMD EPYC CPUs, and 2 TB memory.
The dataset used in this case study is from OpenTI [7], a benchmark traffic dataset featuring initial OD matrices, observed traffic flows, and can generate calibrated OD outputs for varying network conditions. This dataset captures dynamic traffic patterns across different time intervals, making it ideal for evaluating biases and feature quality in generative OD calibration tasks.
@@ 5.8.3 Experimental Design and Methodology
Method Type2: LLM-based O-D Calibration Methods In this section, we have two different designs of LLM-based methods as an exploration, the first one named VanillaLLM is the most straightforward design that, we directly pass the current solution with shape $(T\times 56\times 56)$ and the observed link performance, as well as the calculated gap, and the prompt design is as shown in prompt ①. The second design is based on the combination of the domain expert knowledge and LLM reasoning, which tries to simplify the problem by leveraging heuristic information to reduce the search space and encouraging LLM’s effective reasoning, named ExpertLLM. The difference between the two LLM-related solvers’ structures is as shown in Figure 7.
Figure 7: The difference between the methodology structure of VanillaLLM and ExpertLLM. As shown in the image, the VanillaLLM directly inputs the flattened O-D matrix to the LLM, together with the MSE value, it provides a vast space to optimize given the dimension of 56 x 56 = 3136 for a single time step, and each value is not bounded, this is a very challenging task for LLM to reason. But ExpertLLM first builds the connection from the O-D matrix to link performance and uses human expert’s domain knowledge on the top-k links searching (which causes the observation difference), then it tracks back to the O-D pairs that contribute to these k links, and LLM is asked to solve a problem on a link basis, which is a much easier task for LLM to infer the adjustment direction, this relaxed the difficulty of the proposed question by leveraging necessary information and combine the analysis with reasoning of LLMs.
The following example prompts illustrate the CoT-based OD calibration design:
① Prompt for VanillaLLM solving OD Calibration System Description: We are calibrating a 56×56 Origin-Destination (OD) matrix for a transportation network. Each entry [i, j] represents trips from origin i to destination j. Known Information: 1. Current OD matrix (56 lines): {matrix text} 2. Current MSE: {current MSE} Logics: A lower MSE indicates that the OD matrix better reflects real-world traffic patterns. Task Description: Focus on analyzing patterns and identifying actionable changes to improve matrix accuracy and reduce MSE. Return Constraints: Only reply with 56 lines of a new 56×56 matrix. Example: Adjusted OD Matrix (for 2x2 shape): $\left[\begin{smallmatrix}320&180\\ 160&340\end{smallmatrix}\right]$. Output: Final O-D matrix with error reduction.
@@ 5.8.4 Results and Analysis
- For VanillaLLM, it clearly struggled to optimize the OD Matrix, with MSE values ranging from around 94,000 to well over one million. Despite having full access to the entire OD Matrix and its corresponding MSE, it was unable to derive the actionable insights needed for effective optimization. Intriguingly, the lowest MSE occurred when VanillaLLM generated an all-zero matrix—contrary to explicit instructions—demonstrating its limited capacity to adhere to constraints and reason effectively. This outcome strongly highlights the necessity for human expert intervention to guide and refine the optimization process. Furthermore, our experiments revealed that although VanillaLLM can process extensive data, its internal mechanisms are not well-equipped to handle the complex structure of the OD Matrix, as the model repeatedly defaulted to trivial solutions that minimized error numerically but failed to capture any meaningful spatial or operational nuances. In several cases, the algorithm converged prematurely to an all-zero output, a clear sign of its inability to balance multiple constraints and the multidimensional interplay of variables, resulting in erratic behavior with unpredictable oscillations between high error values and degenerate solutions. This inconsistency suggests that VanillaLLM is overwhelmed by the high dimensionality and vast exploration space inherent in the task, ultimately necessitating a more guided, hybrid approach that incorporates domain-specific heuristics and human oversight to steer the optimization toward robust, contextually aware solutions.
@@
② Prompt for EffectiveLLM solving OD Calibration System Description: We are calibrating a 56x56 Origin-Destination (OD) matrix for a transportation network. Each entry [i, j] represents the number of trips from origin i to destination j. This OD matrix is used in a simulation to generate traffic volume counts on various links in the network. You will be provided with the following details of one link: - Link ID: link id - The simulated volume: simulated vol - The ground truth volume: obs count - The absolute error, which is calculated as: abs(Simulated Volume - Ground Truth Volume) = abs error Known Information: The following {sample size} randomly sampled OD elements (i, j) that contribute to this link, along with their current flow values: {pairs str} Logics: A lower MSE indicates that the OD matrix better reflects real-world traffic patterns. Task Description: Adjust ONLY these OD elements’ flow values to reduce the absolute error, thereby improving the alignment between the simulation results and real-world traffic observations. Return Constraints: - Do not return any placeholder text. - Return ONLY the updated values of the sample size OD elements with their indices, one per line, in the format: [(i, j), new value] [(i2, j2), new value 2] Output: Updated O-D values for selected links.
@@
- For ExpertLLM, as shown in Figure. 8, it achieved significant MSE reductions by focusing on the top K links with the highest absolute error, a strategy that underscores its ability to selectively refine the most critical components of the OD Matrix. This targeted approach not only improves performance by concentrating computational resources on the most problematic links but also demonstrates the model’s potential to deliver rapid improvements in key areas where errors are most obvious. However, despite these promising gains, ExpertLLM eventually reaches a point of diminishing returns where further reductions in MSE plateau, indicating that the automated process alone cannot fully capture the nuanced interdependencies present in the data. At this stage, the benefits of continued unsupervised optimization wane, and the system’s performance becomes constrained by its inability to adapt its focus beyond the initially selected top-K links. In such cases, human intervention becomes crucial to steer the LLM toward more promising avenues for improvement. For instance, replacing the automated top-K selection with input from a transportation planning expert could enable the LLM to identify and prioritize not only the most error-prone links but also those that are strategically important within the overall network, ensuring that improvements have a broader impact on system performance. Moreover, expert guidance might help in reweighting or redefining the criteria for link selection based on real-world constraints and operational priorities, thereby fostering a more context-aware and adaptive optimization process. By integrating domain expertise with automated algorithms, the hybrid approach promises to overcome the plateau in performance, ultimately leading to a more robust and effective optimization of the OD Matrix.
@@ 5.8.5 Summary and Insights
Overall, ExpertLLM demonstrates a significant improvement over both VanillaLLM and the Genetic Algorithm in optimizing the OD Matrix. While the ExpertLLM pipeline shows considerable potential, the performance of both systems ultimately underscores the importance of incorporating human expertise—particularly when improvements plateau—to achieve robust and scalable optimization. Integrating human insight not only enhances the interpretability of LLM suggestions but also provides the necessary domain-specific adjustments to drive further reductions in MSE.
@@
This case study highlights the strengths and limitations of generative AI models in OD calibration tasks:
@@
The results emphasize the importance of integrating intermediate feature generation steps, such as those enabled by Chain-of-Thought reasoning, to enhance OD calibration. Future work will explore methods to further mitigate planning biases and improve model generalizability across diverse traffic networks.
@@ 6.2 Pipelines of Integrating Transportation Planning with Generative AI
  • Spatial and Temporal Correlations: Generative models must capture the intricate spatial-temporal dependencies inherent to transportation systems, such as dynamic OD flows, route interactions, and network bottlenecks. These correlations are further complicated by differences across urban environments, cultural driving habits, and localized traffic patterns. As much of transportation planning is inherently local, results produced by generative models often lack direct transferability. This necessitates careful localization, calibration, and validation to ensure that AI-driven insights align with the specific characteristics of the region being studied. For example, driving behaviors in Los Angeles, New York City, rural areas, or cities in Asia can vary significantly, underscoring the importance of adapting generative AI outputs to reflect these contextual nuances accurately. Without such adaptations, model predictions may fail to meet the practical requirements of local transportation planning.
@@ 6.3 Data Scarcity and the Construction of Domain-Specific Datasets
Data scarcity remains a critical barrier to deploying generative AI in transportation planning. Unlike computer vision or NLP, transportation research lacks large-scale, curated datasets tailored for tasks such as traffic flow prediction, OD calibration, and multimodal optimization [7]. This scarcity arises from high data collection costs, privacy concerns, and the dynamic nature of transportation systems, which require continuous updates. Additionally, the fragmented nature of the transportation planning industry—divided by agency jurisdiction and geographic region—makes it difficult to create standardized, interoperable datasets on a nationwide scale. This lack of coordination limits the availability of shared datasets and hinders the scalability of generative AI applications. Future advancements in data-sharing agreements, standardized formats, and cross-agency collaboration will be essential to overcome these challenges and unlock the full potential of generative AI in transportation planning.
@@
  • OD Calibration: Datasets such as OpenTI [7] integrate OD matrices with observed traffic flows, supporting demand estimation tasks across different traffic environments. However, these datasets often require region-specific adaptations to accommodate local travel behaviors and network structures.
@@
Synthetic Data Generation. Generative AI provides a powerful solution for addressing data scarcity by creating synthetic datasets that replicate real-world transportation patterns. These datasets can fill critical gaps where historical data is unavailable, incomplete, or difficult to collect. For example, generative models can simulate OD matrices for rare events such as natural disasters or large-scale festivals, where traditional data collection methods fall short. Additionally, synthetic data can model mobility patterns in regions with limited transportation records, providing valuable insights for infrastructure planning and policy development.
@@ 6.4 Enhancing Explainability and Reducing Hallucination Risks
Improving Explainability. Ensuring the interpretability of AI-generated transportation insights is critical for decision-making. Generative AI models, particularly those used in OD calibration, travel demand forecasting, and policy simulations, must provide transparent reasoning behind their predictions. Techniques such as feature attribution and causal reasoning can significantly enhance explainability. Feature attribution methods, such as SHAP [209], identify key inputs that influence AI-generated outputs, helping planners understand the relationship between traffic patterns and model predictions. Causal reasoning [292, 293] allows AI models to go beyond correlation-based predictions by explicitly modeling cause-and-effect relationships, improving trust in AI-driven recommendations.
For example, in OD calibration, explainability techniques can highlight how congestion levels, travel demand, and socioeconomic factors contribute to the generated OD matrices, allowing planners to validate whether model outputs align with real-world mobility trends. Similarly, in policy recommendation systems, causal inference can provide insights into how interventions like congestion pricing or transit subsidies impact mode choice and travel behavior, enabling policymakers to evaluate potential trade-offs and unintended consequences before implementation.
@@ 7 Conclusion
  • Domain-specific fine-tuning strategies to enhance performance on tasks like origin-destination (OD) calibration and traffic signal optimization.
@@
  • The creation of benchmark datasets like OpenTI for OD calibration and synthetic traffic analysis, ensuring robust evaluation frameworks for generative AI applications.
An empirical case study on OD flow calibration demonstrates the practical implementation of these methods, illustrating the value of generative AI in improving the efficiency and scalability of complex transportation systems. Finally, we identify critical research directions to address the open challenges in the field:

### v2

@@ 5.8 Case Study: LLM’s Ability in OD-Calibration in Terms of Feature Quality
This subsection presents a case study on how LLMs perform in Origin-Destination (OD) Calibration tasks for transportation planning. Specifically, we utilize an O-D calibration dataset from Open Traffic Intelligence (OpenTI) [7], which includes initial OD matrices and observed traffic counts, to evaluate how generative AI models perform when tasked with O-D calibration. The study focuses on two critical aspects: (1) Exploring the few-shot learning solution designs in current LLMs for solving the O-D calibration task, and (2) assessing the quality of model-generated features (observation counts) compared to ground-truth OD matrix’s count feature, an overview of the problem is shown in Figure 6.
@@ 5.8.1 Task Description
In this section, we consider an origin-destination (O-D) matrix $\mathbf{X}$ of dimension $T\times 56\times 56$, where $T$ is the total number of discrete time slices, and $56$ represents the number of blocks in the city. Specifically, $\mathbf{X}$ contains the entries $x_{t,i,j}$ for $t=1,\dots,T$ and $i,j\in\{1,\dots,56\}$, where $x_{t,i,j}\geq 0$ denotes the traffic demand (number of vehicles) from block $i$ to block $j$ during time slice $t$. Once the full matrix $\mathbf{X}$ is specified, a traffic simulator [280] is executed to generate network flows; however, since the observation point is sparse, the effect from the $f\bigl{(}\mathbf{X}\bigr{)}$ is very limited, which brings much difficulty to this task. We denote $O_{t}$ as the ground-truth measurement at the sensor for time slice $t$. Our goal is to calibrate the unknown entries of $\mathbf{X}$ so as to minimize the mean squared error between the simulator-derived flows and the real observations. Concretely, we define the following:
@@ 5.8.2 Model Configurations, Computational Resources, and Dataset Selection
We conduct exploration on the open-sourced and widely-used generative AI models Llama 3.1-70B [281]—to simulate O-D calibration scenarios. Experiments are conducted on hardware configurations tailored for large-scale traffic data tasks and are deployed on a server with 8 NVIDIA A100 GPUs, 2 AMD EPYC CPUs, and 2 TB memory.
The dataset used in this case study is from OpenTI [7], a benchmark traffic dataset featuring initial O-D matrices, observed traffic flows, and can generate calibrated O-D outputs for varying network conditions. This dataset captures dynamic traffic patterns across different time intervals, making it ideal for evaluating biases and feature quality in generative O-D calibration tasks.
@@ 5.8.3 Experimental Design and Methodology
Method Type2: LLM-based O-D Calibration Methods
In this section, we have two different designs of LLM-based methods as an exploration. The first one, named VanillaLLM, is the most straightforward design that, we directly pass the current solution with shape $(T\times 56\times 56)$ and the observed link performance, as well as the calculated gap, and the prompt design is as shown in prompt ①. The second design is based on the combination of domain expert knowledge and LLM reasoning, which tries to simplify the problem by leveraging heuristic information to reduce the search space and encouraging LLM’s effective reasoning, named ExpertLLM. The difference between the two LLM-related solvers’ structures is as shown in Figure 7.
Figure 7: The difference between the methodology structure of VanillaLLM and ExpertLLM. As shown in the image, the VanillaLLM directly inputs the flattened O-D matrix to the LLM, together with the MSE value, it provides a vast space to optimize given the dimension of 56 x 56 = 3136 for a single time step, and each value is not bounded, this is a very challenging task for LLM to reason. But ExpertLLM first builds the connection from the O-D matrix to link performance and uses human expert’s domain knowledge on the top-k links searching (which reduces the observation space), then it tracks back to the O-D pairs that contribute to these k links, and LLM is asked to solve a problem on a link basis, which is a much easier task for LLM to infer the adjustment direction, this relaxed the difficulty of the proposed question by leveraging necessary information and combine the analysis with reasoning of LLMs.
The following example prompts illustrate the CoT-based O-D calibration design:
① Prompt for VanillaLLM solving O-D Calibration System Description: We are calibrating a 56×56 Origin-Destination (O-D) matrix for a transportation network. Each entry [i, j] represents trips from origin i to destination j. Known Information: 1. Current O-D matrix (56 lines): {matrix text} 2. Current MSE: {current MSE} Logics: A lower MSE indicates that the O-D matrix better reflects real-world traffic patterns. Task Description: Focus on analyzing patterns and identifying actionable changes to improve matrix accuracy and reduce MSE. Return Constraints: Only reply with 56 lines of a new 56×56 matrix. Example: Adjusted O-D Matrix (for 2x2 shape): $\left[\begin{smallmatrix}320&180\\ 160&340\end{smallmatrix}\right]$. Output: Final O-D matrix with error reduction.
@@ 5.8.4 Results and Analysis
- For VanillaLLM, it clearly struggled to optimize the O-D Matrix, with MSE values ranging from around 94,000 to well over one million. Despite having full access to the entire O-D Matrix and its corresponding MSE, it was unable to derive the actionable insights needed for effective optimization. Intriguingly, the lowest MSE occurred when VanillaLLM generated an all-zero matrix—contrary to explicit instructions—demonstrating its limited capacity to adhere to constraints and reason effectively. This outcome strongly highlights the necessity for human expert intervention to guide and refine the optimization process. Furthermore, our experiments revealed that although VanillaLLM can process extensive data, its internal mechanisms are not well-equipped to handle the complex structure of the O-D Matrix, as the model repeatedly defaulted to trivial solutions that minimized error numerically but failed to capture any meaningful spatial or operational nuances. In several cases, the algorithm converged prematurely to an all-zero output, a clear sign of its inability to balance multiple constraints and the multidimensional interplay of variables, resulting in erratic behavior with unpredictable oscillations between high error values and degenerate solutions. This inconsistency suggests that VanillaLLM is overwhelmed by the high dimensionality and vast exploration space inherent in the task, ultimately necessitating a more guided, hybrid approach that incorporates domain-specific heuristics and human oversight to steer the optimization toward robust, contextually aware solutions.
@@
② Prompt for ExpertLLM solving O-D Calibration System Description: We are calibrating a 56x56 Origin-Destination (O-D) matrix for a transportation network. Each entry [i, j] represents the number of trips from origin i to destination j. This O-D matrix is used in a simulation to generate traffic volume counts on various links in the network. You will be provided with the following details of one link: - Link ID: link id - The simulated volume: simulated vol - The ground truth volume: obs count - The absolute error, which is calculated as: abs(Simulated Volume - Ground Truth Volume) = abs error Known Information: The following {sample size} randomly sampled O-D elements (i, j) that contribute to this link, along with their current flow values: {pairs str} Logics: A lower MSE indicates that the O-D matrix better reflects real-world traffic patterns. Task Description: Adjust ONLY these O-D elements’ flow values to reduce the absolute error, thereby improving the alignment between the simulation results and real-world traffic observations. Return Constraints: - Do not return any placeholder text. - Return ONLY the updated values of the sample size O-D elements with their indices, one per line, in the format: [(i, j), new value] [(i2, j2), new value 2] Output: Updated O-D values for selected links.
@@
- For ExpertLLM, as shown in Figure. 8, it achieved significant MSE reductions by focusing on the top K links with the highest absolute error, a strategy that underscores its ability to selectively refine the most critical components of the O-D Matrix. This targeted approach not only improves performance by concentrating computational resources on the most problematic links but also demonstrates the model’s potential to deliver rapid improvements in key areas where errors are most obvious. However, despite these promising gains, ExpertLLM eventually reaches a point of diminishing returns where further reductions in MSE plateau, indicating that the automated process alone cannot fully capture the nuanced interdependencies present in the data. At this stage, the benefits of continued unsupervised optimization wane, and the system’s performance becomes constrained by its inability to adapt its focus beyond the initially selected top-K links. In such cases, human intervention becomes crucial to steer the LLM toward more promising avenues for improvement. For instance, replacing the automated top-K selection with input from a transportation planning expert could enable the LLM to identify and prioritize not only the most error-prone links but also those that are strategically important within the overall network, ensuring that improvements have a broader impact on system performance. Moreover, expert guidance might help in reweighting or redefining the criteria for link selection based on real-world constraints and operational priorities, thereby fostering a more context-aware and adaptive optimization process. By integrating domain expertise with automated algorithms, the hybrid approach promises to overcome the plateau in performance, ultimately leading to a more robust and effective optimization of the O-D Matrix.
@@ 5.8.5 Summary and Insights
Overall, ExpertLLM demonstrates a significant improvement over both VanillaLLM and the Genetic Algorithm in optimizing the O-D Matrix. While the ExpertLLM pipeline shows considerable potential, the performance of both systems ultimately underscores the importance of incorporating human expertise—particularly when improvements plateau—to achieve robust and scalable optimization. Integrating human insight not only enhances the interpretability of LLM suggestions but also provides the necessary domain-specific adjustments to drive further reductions in MSE.
@@
This case study highlights the strengths and limitations of generative AI models in O-D calibration tasks:
@@
The results emphasize the importance of integrating intermediate feature generation steps, such as those enabled by Chain-of-Thought reasoning, to enhance O-D calibration. Future work will explore methods to further mitigate planning biases and improve model generalizability across diverse traffic networks.
@@ 6.2 Pipelines of Integrating Transportation Planning with Generative AI
  • Spatial and Temporal Correlations: Generative models must capture the intricate spatial-temporal dependencies inherent to transportation systems, such as dynamic O-D flows, route interactions, and network bottlenecks. These correlations are further complicated by differences across urban environments, cultural driving habits, and localized traffic patterns. As much of transportation planning is inherently local, results produced by generative models often lack direct transferability. This necessitates careful localization, calibration, and validation to ensure that AI-driven insights align with the specific characteristics of the region being studied. For example, driving behaviors in Los Angeles, New York City, rural areas, or cities in Asia can vary significantly, underscoring the importance of adapting generative AI outputs to reflect these contextual nuances accurately. Without such adaptations, model predictions may fail to meet the practical requirements of local transportation planning.
@@ 6.3 Data Scarcity and the Construction of Domain-Specific Datasets
Data scarcity remains a critical barrier to deploying generative AI in transportation planning. Unlike computer vision or NLP, transportation research lacks large-scale, curated datasets tailored for tasks such as traffic flow prediction, O-D calibration, and multimodal optimization [7]. This scarcity arises from high data collection costs, privacy concerns, and the dynamic nature of transportation systems, which require continuous updates. Additionally, the fragmented nature of the transportation planning industry—divided by agency jurisdiction and geographic region—makes it difficult to create standardized, interoperable datasets on a nationwide scale. This lack of coordination limits the availability of shared datasets and hinders the scalability of generative AI applications. Future advancements in data-sharing agreements, standardized formats, and cross-agency collaboration will be essential to overcome these challenges and unlock the full potential of generative AI in transportation planning.
@@
  • O-D Calibration: Datasets such as OpenTI [7] integrate O-D matrices with observed traffic flows, supporting demand estimation tasks across different traffic environments. However, these datasets often require region-specific adaptations to accommodate local travel behaviors and network structures.
@@
Synthetic Data Generation. Generative AI provides a powerful solution for addressing data scarcity by creating synthetic datasets that replicate real-world transportation patterns. These datasets can fill critical gaps where historical data is unavailable, incomplete, or difficult to collect. For example, generative models can simulate O-D matrices for rare events such as natural disasters or large-scale festivals, where traditional data collection methods fall short. Additionally, synthetic data can model mobility patterns in regions with limited transportation records, providing valuable insights for infrastructure planning and policy development.
@@ 6.4 Enhancing Explainability and Reducing Hallucination Risks
Improving Explainability. Ensuring the interpretability of AI-generated transportation insights is critical for decision-making. Generative AI models, particularly those used in O-D calibration, travel demand forecasting, and policy simulations, must provide transparent reasoning behind their predictions. Techniques such as feature attribution and causal reasoning can significantly enhance explainability. Feature attribution methods, such as SHAP [209], identify key inputs that influence AI-generated outputs, helping planners understand the relationship between traffic patterns and model predictions. Causal reasoning [292, 293] allows AI models to go beyond correlation-based predictions by explicitly modeling cause-and-effect relationships, improving trust in AI-driven recommendations.
For example, in O-D calibration, explainability techniques can highlight how congestion levels, travel demand, and socioeconomic factors contribute to the generated O-D matrices, allowing planners to validate whether model outputs align with real-world mobility trends. Similarly, in policy recommendation systems, causal inference can provide insights into how interventions like congestion pricing or transit subsidies impact mode choice and travel behavior, enabling policymakers to evaluate potential trade-offs and unintended consequences before implementation.
@@ 7 Conclusion
  • Domain-specific fine-tuning strategies to enhance performance on tasks like origin-destination (O-D) calibration and traffic signal optimization.
@@
  • The creation of benchmark datasets like OpenTI for O-D calibration and synthetic traffic analysis, ensuring robust evaluation frameworks for generative AI applications.
An empirical case study on O-D flow calibration demonstrates the practical implementation of these methods, illustrating the value of generative AI in improving the efficiency and scalability of complex transportation systems. Finally, we identify critical research directions to address the open challenges in the field: